OpenStudy (anonymous):

Can someone show me how to solve this? When it was 4 years old, the annual registration fee for a car was $340. When it was 6 years old, the registration fee dropped to $270. If the relationship is linear, write an equation that gives the registration fee f in dollars for the van when it is x years old.

OpenStudy (anonymous):

These are the answers from my hw a. f = -30x - 480 b. f = -35x + 480 c. f = -35x + 470 d. f = -30x + 480 e. f = -30x + 470

OpenStudy (anonymous):

Every 2 years the fee reduces by 340 - 270 = $70 And because it is linear we can assume that every 1 year the fee is reduced by 70/2 = $35 480 f = -35x + 480 Note* that 35 is -35 because every year the fee is decreasing and not increasing. :) And a tip: In a exam if your stuck, just test out each equation to see if it gives you the actual values that are written in the question!

OpenStudy (anonymous):

How would I plug it in to see which one is right to get the values in the question? @javawarrior

OpenStudy (anonymous):

well for example in this question, it gives you the info: 1) When it was 4 years old, the annual registration fee for a car was $340. 2) When it was 6 years old, the registration fee dropped to $270. and if these were the answers: a. f = -30x - 480 b. f = -35x + 480 c. f = -35x + 470 d. f = -30x + 480 e. f = -30x + 470 simply put x as 4 years and see if the equation gives you the same amount as stated in the question. Know what I mean?

OpenStudy (anonymous):

if it gives u the same amount as stated in the question then you know that the equation works and is true! and if it doesnt give you the right amount, you know that the equation is not true!

OpenStudy (anonymous):

Ok I see !! @javawarrior Thanks you are so much help <3 :)